# Scalable MatMul-free Language Modeling

### v5

@@ 3.4 Training Details
To handle non-differentiable functions such as the Sign and Clip functions during backpropagation, we use the straight-through estimator (STE) [43] as a surrogate function for the gradient. STE allows gradients to flow through the network unaffected by these non-differentiable functions, enabling the training of our quantized model. This technique is widely adopted in BNNs and SNNs.
@@
When training a language model with ternary weights, using the same learning rate as regular models can lead to excessively small updates that have no impact on the clipping operation. This prevents weights from being effectively updated and results in biased gradients and update estimates based on the ternary weights. To address this challenge, it is common practice to employ a larger learning rate when training binary or ternary weight language models, as it facilitates faster convergence [44, 45, 11]. In our experiments, we maintain consistent learning rates across both the 370M and 1.3B models, aligning with the approach described in Ref. [46]. Specifically, for the Transformer++ model, we use a learning rate of $43e-$ 4, while for the MatMul-free LM, we employ a learning rate of $34e-$ 3, $32.5e-$ 3, $31.5e-$ 3 in 370M, 1.5B and 2.7B, respectively. These learning rates are chosen based on the most effective hyperparameter sweeps for faster convergence during the training process.
@@
When training conventional Transformers, it is common practice to employ a cosine learning rate scheduler and set a minimal learning rate, typically 0.1×0.1$\times$0.1 × the initial learning rate. We follow this approach when training the full precision Transformer++ model. However, for the MatMul-free LM, the learning dynamics differ from those of conventional Transformer language models, necessitating a different learning strategy. We begin by maintaining the cosine learning rate scheduler and then reduce the learning rate by half midway through the training process. Interestingly, we observed that during the final training stage, when the network’s learning rate approaches 0, the loss decreases significantly, exhibiting an S-shaped loss curve. This phenomenon has also been reported by  [11, 44] when training binary/ternary language models.
4 Experiments
Our primary focus is testing the MatMul-free LM on moderate-scale language modeling tasks. We compare two variants of our MatMul-free LM against a reproduced advanced Transformer architecture (Transformer++, based on Llama-2) across three model sizes: 370M, 1.3B, and 2.7B parameters. For a fair comparison, all models are pre-trained on the SlimPajama dataset [47], with the 370M model trained on 15 billion tokens, and the 1.3B and 2.7B models trained on 100 billion tokens each. All experiments were conducted using the flash-linear-attention [48] framework, with the Mistral [42] tokenizer (vocab size: 32k) and optimized triton kernel. The training of our models was conducted using 8 NVIDIA H100 GPUs. The training duration was approximately 5 hours for the 370M model, 84 hours for the 1.3B model, and 173 hours for the 2.7B model.
4.1 Scaling Law of MatMul-free LM
Neural scaling laws posit that model error decreases as a power function of training set size and model size, and have given confidence in performance. Such projections become important as training becomes increasingly expensive with larger models. A widely adopted best practice in LLM training is to first test scalability with smaller models, where scaling laws begin to take effect [49, 50, 51]. The GPT-4 technical report revealed that a prediction model just $1/10,000$ the size of the final model can still accurately forecast the full-sized model performance [52].
4.2 Downstream Tasks
In line with benchmarking in BitNet, we evaluated the zero-shot performance of these models on a range of language tasks, including ARC-Easy [53], ARC-Challenge [53], Hellaswag [54], Winogrande [55], PIQA [56], and OpenbookQA [57]. The results are shown in Tab. 1. Details about the datasets can be found in Appendix C. All evaluations are performed using the LM evaluation harness [58]. The MatMul-free LM models achieve competitive performance compared to the Transformer++ baselines across all tasks, demonstrating its effectiveness in zero-shot learning despite the absence of MatMul operations, and the lower memory required from ternary weights. Notably, the 2.7B MatMul-free LM model outperforms its Transformer++ counterpart on ARC-Challenge and OpenbookQA, while maintaining comparable performance on the other tasks. As the model size increases, the performance gap between MatMul-free LM and Transformer++ narrows, which is consistent with the scaling law. These results highlight that MatMul-free architectures are capable achieving strong zero-shot performance on a diverse set of language tasks, ranging from question answering and commonsense reasoning to physical understanding.
Models Size ARCe ARCc HS OQ PQ WGe Avg.
370M parameters with 15B training tokens, Layer=24, d=1024
Transformer++ 370M 45.0 24.0 34.3 29.2 64.0 49.9 41.1
MatMul-free RWKV-4 370M 44.7 22.8 31.6 27.8 63.0 50.3 40.0
Ours 370M 42.6 23.8 32.8 28.4 63.0 49.2 40.3
1.3B parameters with 100B training tokens, Layer=24, d=2048
Transformer++ 1.3B 54.1 27.1 49.3 32.4 70.3 54.9 48.0
MatMul-free RWKV-4 1.3B 52.4 25.6 45.1 31.0 68.2 50.5 45.5
Ours 1.3B 54.0 25.9 44.9 31.4 68.4 52.4 46.2
2.7B parameters with 100B training tokens, Layer=32, d=2560
Transformer++ 2.7B 59.7 27.4 54.2 34.4 72.5 56.2 50.7
Ours 2.7B 58.5 29.7 52.3 35.4 71.1 52.1 49.9
Figure 4: Performance comparison and analysis of different models and configurations. (a) and (b) show the training performance comparison between Vanilla BitLinear and Fused BitLinear in terms of time and memory consumption as a function of batch size. (c) illustrates the effect of learning rate on training loss for the MatMul-free LM. (d) compares the inference memory consumption and latency between MatMul-free LM and Transformer++ across various model sizes.
4.3 Learning Rate Analysis
The learning rate is a crucial hyper-parameter in language model training, and models become more sensitive to the learning rate in the ternary/binary weight regime. To determine the optimal learning rate, we conducted a search within the range of $31.5e-$ 3 to $23e-$ 2 using our 370M model with a batch size of 50k tokens. The results of this search are shown in Fig. 4(c). The results demonstrate that the final training loss monotonically decreases as the learning rate increases from $31.5e-$ 3 to $21e-$ 2. The model only exhibits instability when the learning rate exceeds $22e-$ 2. This finding suggests that previous works employing ternary weights, such as BitNet, which uses a learning rate of $31.5e-$ 3, may not be optimal and that higher learning rates could potentially lead to better performance. These findings align with the observations from the Deepseek LLM [59] which found that the optimal learning rate for conventional LLMs is actually larger than the values typically reported in most LLM training setups. Interestingly, we also observed that models trained with larger learning rates at the start of the training process exhibit a more rapid decrease in training loss during the later stages of training compared to those trained with smaller learning rates.
4.4 Training Efficiency Comparison
Our experiments show that our fused operator benefits from larger batch sizes in terms of faster training speeds and reduced memory consumption. When the batch size is $2^{8}$, the training speed of the 1.3B parameter model improves from 1.52s to 1.21s per iteration, a 25.6% speedup over the vanilla implementation. Additionally, memory consumption decreases from 82GB to 32GB, a 61.0% reduction in memory usage. The performance of the Fused implementation improves significantly with larger batch sizes, allowing more samples to be processed simultaneously and reducing the total number of iterations.
4.5 Inference Efficiency Comparison
Fig. 4(d) presents a comparison of GPU inference memory consumption and latency between the proposed MatMul-free LM and Transformer++ for various model sizes. In the MatMul-free LM, we employ BitBLAS [60] for acceleration to further improve efficiency. The evaluation is conducted with a batch size of 1 and a sequence length of 2048. The MatMul-free LM consistently demonstrates lower memory usage and latency compared to Transformer++ across all model sizes. For a single layer, the MatMul-free LM requires only 0.12 GB of GPU memory and achieves a latency of 3.79 ms, while Transformer++ consumes 0.21 GB of memory and has a latency of 13.87 ms. As the model size increases, the memory and latency advantages of the MatMul-free LM become more pronounced. It is worth noting that for model sizes larger than 2.7B, the results are simulated using randomly initialized weights. For the largest model size of 13B parameters, the MatMul-free LM uses only 4.19 GB of GPU memory and has a latency of 695.48 ms, whereas Transformer++ requires 48.50 GB of memory and exhibits a latency of 3183.10 ms. These results highlight the efficiency gains achieved by the MatMul-free LM, making it a promising approach for large-scale language modeling tasks, particularly during inference.
5 FPGA Implementation and Results
5.1 Implementation
@@ 5.2 Results
By instead using the full 512-bit DDR4 interface and parallelizing the TMATMUL functional unit, which dominates 99% of core processing time, a further speed-up of approximately 64$\times$ is projected, while maintaining the same clock rate without additional optimizations or pipelining, as shown in Table 3. Given the 370M parameter model where $L=24$, $d=512$, the total projected runtime is 16.08ms, and a throughput of approximately 62 tokens per second. The 1.3B parameter model, where $L=24$ and $d=2048$, has a projected runtime of 42ms, and a throughput of 23.8 tokens per second. This reaches human reading speed at an efficiency that is on par with the power consumption of the human brain. This is for the case of a single core with a single batch of data, and can be scaled up significantly through batch processing by pipelining the single core with a negligible increase in average power, or alternatively, by increasing the core count with an increase in power (Table 2).
@@
6 Conclusion
We have demonstrated the feasibility and effectiveness of the first scalable MatMul-free language model. Our work challenges the paradigm that MatMul operations are indispensable for building high-performing language models and paves the way for the development of more efficient and hardware-friendly architectures. We achieve performance on par with state-of-the-art Transformers while eliminating the need for MatMul operations, with an optimized implementation that significantly enhances both training and inference efficiency, reducing both memory usage and latency. As the demand for deploying language models on various platforms grows, MatMul-free LMs present a promising direction for creating models that are both effective and resource-efficient. However, one limitation of our work is that the MatMul-free LM has not been tested on extremely large-scale models (e.g., 100B+ parameters) due to computational constraints. This work serves as a call to action for institutions and organizations that have the resources to build the largest language models to invest in accelerating lightweight models. By prioritizing the development and deployment of MatMul-free architectures such as this one, the future of LLMs will only become more accessible, efficient, and sustainable.
APPENDIX
Appendix A Quantization for MatMul-free Dense Layers
During training, we first quantized the weights to $\{-1,0,1\}$ by using an absmean quantization function, which scales the weight matrix by its average absolute value and rounds each element to the nearest ternary integer among $\{-1,0,+1\}$:
$\widetilde{\mathbf{W}}\in\mathbb{R}^{n\times m}=\mathrm{RoundClip}(\frac{W}{\gamma+\epsilon},-1,1),$
$\mathrm{RoundClip}(x,a,b)=\max(a,\min(b,\mathrm{round}(x))),$
$\gamma=\frac{1}{nm}\sum_{ij}|W_{ij}|,$
where $n$ and $m$ are the number of rows and columns of $W$. After weight quantization, activations are also quantized to 8-bit precision, as is done with BitNet. We use absmax quantization, which scales activations into the range $[-Q_{b},Q_{b}]$, given that $b$ is the number of bits and $Q_{b}=2^{b-1}$:
$\widetilde{x}=\mathrm{Quant}(x)=\mathrm{Clip}(x\times\frac{Q_{b}}{\gamma},-Q_{b}+\epsilon,Q_{b}-\epsilon),$ (6)
$\mathrm{Clip}(x,a,b)=\max(a,\min(b,x)),\quad\gamma=||x||_{\infty}.$ (7)
where $\epsilon$ is a small number that prevents overflow during clipping.
With these quantization equations, the MatMul can be written as:
$y=\widetilde{x}\circledast\widetilde{\mathbf{W}}$
To preserve variance and maintain numerical stability after quantization, we use RMSNorm [61] before activation quantization, which is also used in BitNet:
$y=\widetilde{x}\circledast\widetilde{\mathbf{W}}=~{}\mathrm{Quant}(\mathrm{RMSNorm}(x))\circledast\widetilde{\mathbf{W}}\times\frac{\beta\gamma}{Q_{b}},$ (8)
$\mathrm{RMSNorm}(x)=\frac{x}{\sqrt{\mathrm{E}(x^{2})+\epsilon}},\quad\beta=\frac{1}{nm}||W||_{1},\quad\gamma=||x||_{\infty}.$
where $Q_{b}$ is the max value for activation, and $\beta$ is the mean of the weight matrix.
Appendix B RWKV-4 as a MatMul-free Token Mixer
RWKV-4 can also function as a token mixer which utilizes recurrence to mix temporal information and a 1-D hidden states that is updated using element-wise Hadamard products which avoids MatMul operations. This approach offers several advantages over conventional transformers, including computational efficiency, effective propagation of information across time steps, simplified model architecture, and reduced memory usage. Given the good performance of RWKV-4 in capturing dependencies and relationships between tokens across long-ranges of time steps, we additionally tested a ternary version of RWKV-4, though it underperformed compared to what we proposed in the main manuscript. In the interest of saving the research community compute-hours, we explain the process and report our ‘negative’ results here. The RWKV-4 token mixer can be expressed as follows:
$\mathbf{r}_{t}$ $=(\mathbf{\mu}_{r}\mathbf{x}_{t}+(1-\mathbf{\mu}_{r})\mathbf{x}_{t-1})\circledast\mathbf{W}_{r}\in\mathbb{R}^{1\times d},$
$\mathbf{k}_{t}$ $=(\mathbf{\mu}_{k}\mathbf{x}_{t}+(1-\mathbf{\mu}_{k})\mathbf{x}_{t-1})\circledast\mathbf{W}_{k}\in\mathbb{R}^{1\times d},$
$\mathbf{v}_{t}$ $=(\mathbf{\mu}_{v}\mathbf{x}_{t}+(1-\mathbf{\mu}_{v})\mathbf{x}_{t-1})\circledast\mathbf{W}_{v}\in\mathbb{R}^{1\times d},$
$\mathbf{h}_{t}$ $=\frac{\mathbf{a}_{t-1}+e^{\mathbf{m}+\mathbf{k}_{t}}\odot\mathbf{v}_{t}}{\mathbf{b}_{t-1}+e^{\mathbf{m}+\mathbf{k}_{t}}}\in\mathbb{R}^{1\times d},$
$\mathbf{a}_{t}$ $=e^{-\mathbf{w}}\odot\mathbf{a}_{t-1}+e^{\mathbf{k}_{t}}\odot\mathbf{v}_{t},$
$\mathbf{b}_{t}$ $=e^{-\mathbf{w}}\odot\mathbf{b}_{t-1}+e^{\mathbf{k}_{t}}\in\mathbb{R}^{1\times d},$
$\mathbf{o}_{t}$ $=(\sigma(\mathbf{r}_{t})\odot\mathbf{h}_{t})\circledast\mathbf{W}_{o}\in\mathbb{R}^{1\times d},$
$\mathbf{a}_{0}$ $=0\in\mathbb{R}^{1\times d},$
$\mathbf{b}_{0}$ $=0\in\mathbb{R}^{1\times d},$

### v6

@@ 6.4 Training Details
To handle non-differentiable functions such as the Sign and Clip functions during backpropagation, we use the straight-through estimator (STE) ste as a surrogate function for the gradient. STE allows gradients to flow through the network unaffected by these non-differentiable functions, enabling the training of our quantized model. This technique is widely adopted in BNNs and SNNs.
@@
When training a language model with ternary weights, using the same learning rate as regular models can lead to excessively small updates that have no impact on the clipping operation. This prevents weights from being effectively updated and results in biased gradients and update estimates based on the ternary weights. To address this challenge, it is common practice to employ a larger learning rate when training binary or ternary weight language models, as it facilitates faster convergence zhang2024binarized ; liu2023binary ; ma20241bitllm . In our experiments, we maintain consistent learning rates across both the 370M and 1.3B models, aligning with the approach described in Ref. qin2023scaling . Specifically, for the Transformer++ model, we use a learning rate of $43e-$ 4, while for the MatMul-free LM, we employ a learning rate of $34e-$ 3, $32.5e-$ 3, $31.5e-$ 3 in 370M, 1.5B and 2.7B, respectively. These learning rates are chosen based on the most effective hyperparameter sweeps for faster convergence during the training process.
@@
When training conventional Transformers, it is common practice to employ a cosine learning rate scheduler and set a minimal learning rate, typically 0.1×0.1$\times$0.1 × the initial learning rate. We follow this approach when training the full precision Transformer++ model. However, for the MatMul-free LM, the learning dynamics differ from those of conventional Transformer language models, necessitating a different learning strategy. We begin by maintaining the cosine learning rate scheduler and then reduce the learning rate by half midway through the training process. Interestingly, we observed that during the final training stage, when the network’s learning rate approaches 0, the loss decreases significantly, exhibiting an S-shaped loss curve. This phenomenon has also been reported by  ma20241bitllm ; zhang2024binarized when training binary/ternary language models.
6.5 Implementation Details on Intel Loihi 2
In order to run the MatMul-free LM on the Intel Loihi 2 chip, the original model must be quantized, implemented using fixed-point arithmetic using custom microcode on the Loihi 2 platform, and finally verified against the original model to make sure that the mismatch between the fixed-point model on Loihi 2 and the original model on GPU is limited. We present these steps in the following sections.
Model Quantization
As a first step, we quantize all weights and activations of the MatMul-free LM and verify the resulting performance of the quantized model on the downstream tasks presented previously. All weight matrices are already ternarized, thus not needing further quantization. However, every BitLinear layer with a ternary weight matrix $W\in\mathbb{R}^{d\times m}$ is preceded by a root mean square normalization (RMSNorm) operation, given by:
$\text{RMSNorm}(x;w,\epsilon)=\frac{x}{\sqrt{\epsilon+\sum_{i}^{d}x_{i}^{2}}}\odot w$ (6)
[…]
Whereas activations of the MatMul-free LM are stored as 16-bit floating-point numbers, activations on Loihi 2 are sent between neurons as integer payloads. To test the performance degradation of the model when quantization activations to lower bit precision, the model was quantized using dynamic quantization in PyTorch, where quantization scales for the activations $s_{a}$ are computed for each batch individually. When quantizing activations in the RMSNorm layer, the $\epsilon$ value underflows in most layers, potentially leading to division by zero. We therefore set $\epsilon=10^{-3}$ and report results with this modified value. Table 2 shows that for the 370M model, performance drops by less than 1.5% relative to the original model when using 16-bit activation quantization. Performance drops more significantly, by over 12% relative to the original model, when using 8-bit activations. All further results will assume 8-bit weights and 16-bit activations. We note that these results are for naive post-training quantization and may be significantly improved through quantization-aware training or fine-tuning the model with specific quantization requirements.
Fixed-Point Arithmetic
Once all quantization scales are computed, the model can be run in static quantization with integer weights and activations. However, some operations used in the MatMul-free LM are not defined on integers, namely the sigmoid activation function $\sigma$ and the inverse-square-root in the RMSNorm.
We employ a look-up-table (LUT) for a fixed-point approximation of the logistic sigmoid function, $\sigma(x)=1/(1+e^{-x})$. Specifically, we scale the floating-point input $x$ by $2^{x_{\text{exp}}}$ where $x_{\text{exp}}=6$, quantize it to an integer domain, and store precomputed $\lfloor\sigma\!\bigl{(}\frac{x}{2^{x_{\text{exp}}}}\bigr{)}\cdot 2^{y_{\text{exp}}}\rfloor$ values in a LUT for positive inputs. For negative inputs, we exploit $\sigma(-x)=1-\sigma(x)$, thus requiring only half-range values. During inference, a piecewise linear interpolation between adjacent LUT entries refines the output. This design offers efficient computation and controllable approximation error.
For the inverse-square-root in the RMSNorm layer, we adapted a well-known “fast inverse square root” algorithm Fast InvSqrt to operate entirely in fixed-point arithmetic. Our method treats the input $\tilde{x}$ as an integer paired with a fixed exponent, and uses a LUT with 24 values to produce an initial guess for $\sqrt{\tilde{x}}$. This estimate is then refined using five iterations of the Newton-Raphson method, all in a fixed-point format. See Appendix C.2 for more details.
Mapping the MatMul-free LM onto Loihi
The model on Loihi 2 is implemented as a set of neurons that are connected through synapses. Each neuron asynchronously executes a simple microcode program and sends its output to a collection of other neurons through synapses. Because every neuron contains only information pertaining to its own dynamics, aggregate operations–such as the sum of squares over an activation vectors–must be implemented through an additional neuron that receives inputs from all neurons in the relevant layer. Fig. 9 in Appendix C.4.2 shows the partitioning of the MatMul-free model into Loihi-compatible neurons.
Appendix A FPGA Implementation and Results
Neither GPUs nor Intel Loihi 2 support 2-bit weights, so we additionally did an RTL-only implementation of the MatMul-free LM on a D5005 Stratix 10.
A.1 Implementation
@@ A.2 Results
By instead using the full 512-bit DDR4 interface and parallelizing the TMATMUL functional unit, which dominates 99% of core processing time, a further speed-up of approximately 64$\times$ is projected, while maintaining the same clock rate without additional optimizations or pipelining, as shown in Table 5. Given the 370M parameter model where $L=24$, $d=512$, the total projected runtime is 16.08ms, and a throughput of approximately 62 tokens per second. The 1.3B parameter model, where $L=24$ and $d=2048$, has a projected runtime of 42ms, and a throughput of 23.8 tokens per second. Despite not being optimized for 2-b precision weights, Intel Loihi 2 results outperform the FPGA implementation in terms of both throughput and power likely due to a mix of efficient packet routing, deep optimizations for recurrent state updates, unused digital signal processing (DSP) units on the FPGA, and process node deviations. This is for the case of a single core with a single batch of data, and can be scaled up significantly through batch processing by pipelining the single core with a negligible increase in average power, or alternatively, by increasing the core count with an increase in power (Table 4).
@@
Appendix B Proof of Linear Scaling in the Proposed MatMul-free Linear Gated Recurrent Unit
The MatMul-free Linear Gated Recurrent Unit (MLGRU) presents a more efficient structure compared to the conventional Key-Value (KV) Cache mechanism used in Transformer models. To demonstrate its advantages, we examine both models in terms of computational complexity and memory requirements with respect to the sequence length $T$.
In Transformers, the KV Cache mechanism stores all keys and values from previous time steps for autoregressive generation. At each time step $t$, the model requires three steps: computing the query, key, and value vectors for the input $\boldsymbol{x}_{t}$, followed by calculating the attention weights by comparing the query $\boldsymbol{q}_{t}$ with all previous keys, and finally constructing the context vector by a weighted sum of the previous values. Formally, the query, key, and value are calculated as
$\boldsymbol{q}_{t}=\boldsymbol{x}_{t}\mathbf{W}_{q},\quad\boldsymbol{k}_{t}=\boldsymbol{x}_{t}\mathbf{W}_{k},\quad\boldsymbol{v}_{t}=\boldsymbol{x}_{t}\mathbf{W}_{v},$
where $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{attn}}}$ are weight matrices. The attention weights, calculated as
$\alpha_{ti}=\frac{\exp\left(\frac{\boldsymbol{q}_{t}^{\top}\boldsymbol{k}_{i}}{\sqrt{d_{\text{attn}}}}\right)}{\sum_{j=1}^{t}\exp\left(\frac{\boldsymbol{q}_{t}^{\top}\boldsymbol{k}_{j}}{\sqrt{d_{\text{attn}}}}\right)},\quad\text{for }i=1,\dots,t,$
lead to the context vector
$\boldsymbol{z}_{t}=\sum_{i=1}^{t}\alpha_{ti}\boldsymbol{v}_{i}.$
These operations grow linearly in complexity with $t$, producing $O(t\cdot d_{\text{attn}})$ computational steps per time step.
The computational complexity of the KV Cache over a sequence of length $T$ therefore follows a quadratic scaling pattern:
$\sum_{t=1}^{T}O(t\cdot d_{\text{attn}})=O(d_{\text{attn}}\cdot T^{2}).$
The memory requirement, dictated by the storage of all previous keys and values, also scales linearly with $T$, resulting in $O(T\cdot d_{\text{attn}})$.
In contrast, the proposed MLGRU architecture achieves simplicity and efficiency by eliminating matrix multiplications and relying on element-wise ternary-weight operations, making it a MatMul-free structure. This architecture uses a recurrent update rule for the hidden state $\boldsymbol{h}_{t}$, leveraging the previous hidden state $\boldsymbol{h}_{t-1}$ and the current input $\boldsymbol{x}_{t}$ without retaining past sequence information. The key equations include the forget gate
$\boldsymbol{f}_{t}=\sigma\left(\boldsymbol{x}_{t}\circledast\mathbf{W}_{f}+\boldsymbol{b}_{f}\right),$